The pitch of a screw is the distance between it's threads. With each complete rotation, the screw goes in or out a distance equal to it's pitch. After 5 complete rotations, the screw is extended 2/12 inch into the wallboard. Find the pitch of the screw.
Answer by Alan3354(69443)

The pitch of a screw is the distance between it's [sic] threads. With each complete rotation, the screw goes in or out a distance equal to it's [sic] pitch. After 5 complete rotations, the screw is extended 2/12 inch into the wallboard. Find the pitch of the screw.
-----------------
2/12 = 1/6 inch
pitch = (1/6)/5 = 1/30 inch
--> 30 TPI (threads per inch)
------------
PS it's = it is